# Review on the decoding algorithms for surface codes

### v4

@@ 6 Discussion
Following the present discussion, and as it has been seen throughout the text, the decoding stage stands as a pivotal operation within the successful functioning of, not only the surface code, but any correcting code in general. One should be aware of the fact that the named code threshold depends on three aspects: the code family in consideration, the decoding method for the code and the noise model in consideration [90, 91]. In this sense, it is straightforward to see that designing decoding algorithms is crucial in terms of the ability of a family of codes to correct errors. The intrinsic locality and degeneracy of the surface code allows for several decoding processes to present themselves as valid candidates for future real device implementation. Nevertheless, in addition to the code threshold for a decoding method, two other critical aspects stand as limiting factors for the implementation of surface codes decoding in a real-time fashion when implemented on experimental hardware: run time and circuit-level noise.
@@
Decoding runtime refers to the actual time that a decoder requires to ouptut an estimate of the error that has corrupted the logical qubit encoded by a code. Hence, this inherently imposes a delay in the error correcting system that is hazardous for its actual operation. This comes from the fact that, once the syndrome measurement has been done, the data qubits will still continue to suffer from errors as they will not stop to decohere. Thus, if the delay between measuring the checks and applying the recovery operation is too high, then the algorithm will fail, with high probability, the actual error configuration when recovery is executed (See appendix A for a discussion of why real-time decoding is required). This is especially important by taking into account that the decoherence times of state-of-the-art qubits are short and the generation of syndrome data is done a fast rate [^31] Note that this is not true for all qubit technologies such as ion traps, but those present additional problems such as the fact that their operation times, quantum gates, are very slow. [30]. As seen in through this tutorial, it is pretty fair to state that maintaining code threshold while reducing the actual complexity of a decoder is a hard problem. This is a result of the trade-off of being fast against considering more error configurations, which actually relates to code performance. This logic comes from the fact that the traditional comparison of accurate and fast is done in terms of code threshold and worst case complexity. However, it is a recent trend to benchmark runtime by expected complexity, i.e. an average of the number of operation required to decode different error patterns. The rationale behind this is that worst-case events are generally exponentially rare, implying that their impact will not be as high in code performance as other more frequent cases. This effect should be more important whenever hardware noise improves lower to sub-threshold physical error rates and, thus, expected runtime might be a better benchmark for decoding algorithms. In this sense, designing faster decoding algorithms that are able to maintain the capabilities of the code in terms of decoding accuracy is a critical issue for the QEC field. It is important to state that significant advances in this direction are being obtained with proposals such as Sparse Blossom [57] or Fussion Blossom [119]. Furthermore, recently, sliding-window decoding schemes have been proposed that allow parallel decoding of the syndrome data by means of matching like decoding algorithms (MWPM and UF), thus, dealing with the exponential backlog problem of real-time decoding while maintaining logical fidelity [227, 228].
@@
  • The rectangular surface code: using a surface code with rectangular shape makes the minimal weight logical error $Z_{L}$ to require more physical $Z$ operators, increasing the $Z$-distance $d_{Z}$ and, consequently, increasing the number of $Z$-errors it can correct: $(d_{Z}-1)/2$ [131]. On the negative side, this is at the expense of increasing the number of necessary data qubits to $d_{X}d_{Z}$. In FIG. 28, a rectangular code of distances $d_{X}=5,d_{Z}=7$ is shown, notice how the logical operator $Z_{L}$ requires additional Pauli operators than for the $X_{L}$ operator. This code construction may use the reviewed decoders in order to estimate the channel errors.
@@
    Figure 28: Graphical representation of a $d_{X}=5,d_{Z}=7$ rectangular rotated planar code. $X_{L}$ and $Z_{L}$ operators are also represented.
@@
  • Clifford deformation: in Clifford deformed surface codes, the check qubits are conjugated respect to a unitary operator from the Clifford group [230] so that other valid stabilizers are obtained [106]. By means of this stabilizer modifications, the new stabilizer codes can be constructed so that they are more susceptible to $Z$-noise. For example, the XY-code is a surface code which, instead of having $X$- and $Z$-checks, has $Y$- and $X$-checks, both of which are susceptible to $Z$-noise. A graphical depiction of the XY-code is given in the top figure of FIG. 29. As a result of such susceptibility to $Z$-errors, the associated decoder can make use of the correlation between $X$ and $Y$-errors, i.e. $XY=Z$, in order to be more accurate in detecting $Z$-noise [170, 87, 88]. Another popular Clifford deformed surface code is the XZZX surface code [41, 231]. For this code, all check qubits act equally on their surrounding data qubits operating two $X$ and two $Z$ operators as shown in the bottom figure of FIG. 29. One can see that the commutation between the checks is maintained, since adjacent checks either anti-commute twice or directly commute. An interesting characteristic of the XZZX code is that there is only one set of $Z$ physical operators which produce a logical operator. As is shown in FIG. 29, all the products of this logical operator with the stabilizer set produce physical operators which are composed of $X$, $Y$ and $Z$-operators. This feature complicates the probability of a logical phase-flip, $Z_{L}$, error under a highly $Z$-biased noise model. Moreover, for the case in which the $Z$-bias is infinite, e.g. a pure dephasing channel, the XZZX surface code operates as a series of disjoint repetition codes that can be decoded independently. The data qubits of such repetition codes are the top-left to bottom-right diagonals of the overall XZZX code, as can be seen in the bottom of FIG. 29, where the diagonal consisting of two $Z$ physical errors produces syndromes in their top-left and bottom-right sides. Given this condition, the threshold of the XZZX code reaches an outstanding $50\%$ [41, 231], while reducing the amount of required qubits significantly. Therefore, it can be seen that a bias towards $Z$-errors is not only not problematic, but also beneficial if the code and decoder are tailored for such model.
@@
    Figure 29: On the top, a graphical representation of a rotated planar XY code with the operators of the stabilizer generators. On the bottom, the same representation but for the rotated planar XZZX code altogether with a $Z_{L}$ operator
@@
  • $X/Z$ correlation: another important approach to improve the performance of surface codes is to make decoders aware of the fact that $Y$-errors represent a correlation between the $X$ and $Z$ errors. One downside of the independent $X$ and $Z$ decoding of CSS codes such as the surface code is that it makes the decoding blind of the correlation among those errors, i.e. it is usually considered that $p_{Y}=p_{X}p_{Z}$. However, as seen in Section 4, the Pauli noise models have $p_{X}\approx p_{Y}$ (both in biased and depolarizing) and, thus, by not taking this into account the decoder underestimates the frequency in which $Y$-errors corrupt the data qubits of the code. A method in order to circumvent this problem (without explicitly changing the structure of the code) consists in first decoding for one of the errors and, afterwards, using the partial recovered error in order to decode the other error. E.g., given an error in a CSS surface code, first decode the $X$-error, and then consider that, for recovering for the $Z$-error, every qubit has an updated probability of undergoing a $Z$-error defined by:
@@ Acknowledgements
We want to acknowledge Nicolas Delfosse, Pavel Panteleev, Christopher Chubb, David Tuckett, Michael Newman, Manabu Hagiwara, Iñigo Barasoain and Javier Oliva for fruitful discussions.
@@ Appendix A The backlog problem and why QEC requires real-time decoding
As discussed in the main text, the decoding algorithms for the surface code need to be fast enough to operate in real time. This is necessary because running fault-tolerant algorithms outside the reach of classical simulation requires the use of quantum algorithms that apply $T=\text{diag}(1,e^{i\frac{\pi}{4}})$ gates over the logical space of the code. The use of $T$ gates guarantees that the code supports the universal gate set Clifford+T that can be used to compile any quantum computation, a result known from the Solovay-Kitaev theorem [237]. However, the Eastin-Knill theorem states that no quantum error correction code can implement a universal gate set in a transversal manner  [^32] A transversal logical gate is one that is implemented by performing physical gates over the physical qubits of the code in a manner that avoids the propagation of physical errors throughout the code. For example, the surface code allows the application of a logical $X$-gate by applying $X$ operators on the physical qubits horizontally from boundary to boundary (Recall figure 4).. In this regard, the surface code supports transversal implementation of logical Clifford gates but lacks a transversal version of the logical $T$-gate. Fortunately, there are ways to circumvent the Eastin-Knill theorem with the purpose of achieving universal fault-tolerant computation. The most popular method is known as magic state injection [227]. A magic state injection protocol can be seen in FIG. 30, where the protocol requires the distillation of a magic state [56] and the injection of such a state to the logical qubit by means of a teleportation circuit.
@@
Coming back to the topic of decoding algorithms, note that if one restricts the operation of the QEC to the set of Clifford gates, then due to the fact that the Clifford group is the normalizer of the Pauli group, it is not necessary to know the logical state for applying the circuit [56]. Since Clifford gates map Paulis into Paulis, the only thing that changes is the interpretation of the final measurements after the computation. Therefore, when running such circuits, one can record the syndrome history during the computation and post-process the errors using the recorded syndrome information. In this sense, one can construct a quantum memory without requiring real-time decoding [56]. However, the scenario drastically changes whenever useful computations are taken into account [^33] By useful we mean quantum computations that cannot be efficiently simulated by classical means and, hence, require $T$-gates.. That is because when injecting the magic state, one must be certain if the $S$-gate after the measurement should be applied or not. Such a decision can only be reliably performed if the syndrome data before the application of said gate is decoded [227]. If the correction of such a Clifford error is ignored, the resulting error might propagate into a much more complicated error that will corrupt the computation [56]. Therefore, in a fault-tolerant quantum computer, the classical processing of the syndrome data, that is, the decoding algorithm; should be fast enough so that syndrome data does not accumulate into an undecodable degree. This accumulation of syndrome data due to the latency of the decoding algorithm is usually referred as the backlog problem [56, 227]. $r_{\mathrm{proc}}$ is defined as the rate at which the decoder processes the syndrome data while $r_{\mathrm{gen}}$ is the rate at which the syndrome data is generated (both quantities in bauds). As discussed by Terhal in [56], if the ratio $r_{\mathrm{gen}}/r_{\mathrm{proc}}=f>1$, then a small initial backlog in the processing of the syndrome data will result in an exponential slow down as a function of the $T$-gate depth of the algorithm. More concretely, Terhal lower bounded the running time of $k$ $T$-depth quantum algorithms as $cf^{k}$ when $f>1$, for some constant $c$ [56, 227]. Thus, the straightforward conclusion is that decoding algorithms must be fast enough to ensure that the injections of the magic states are delayed by only a constant amount of time (instead of an exponentially increasing amount of time).
@@ Appendix B Multiple-round decoding for measurement errors
Throughout this article, surface code decoders have been described as correcting errors arising from data qubit errors, i.e. considering that check measuremnts are noiseless. However, this simplified perspective is overly optimistic because circuit-level noise arising from gate and SPAM errors make syndrome measurement to be noisy too. As mentioned in the main article section on error models and decoding, circuit level noise will lead to incorrect syndromes. If we attempt to decode an erroneous syndrome, the recovered error will likely result in a logical error. Therefore, a new approach is needed to account this type of errors in the code. The common technique consists in performing multiple syndrome extracting rounds. When a check qubit undergoes a change in its measurement value, it implies a non-trivial syndrome element for such measurement round. Subsequently, decoding algorithms can be employed by connecting check qubits within a particular time instance to themselves in the following time, creating a graph containing space and time information of the errors. Recall that not every decoding algorithm requires this multiple-round measurements as, for example, the cellular automaton decoders [124, 186]. These type of decoders are usually known as single-shot decoders [161]. It is important to say that not every QEC family can admit a single-shot decoding algorithm [161].
Figure 31: Graphical representation of an error in a space-time surface code.
@@
The common approach for considering measurement error noise channels consists in introducing an error probability to the state initialization of the measurement qubit, noisy Haddamard and CNOT gates as described in the error model section and noisy measurement probabilities. Often, this is done by considering all gates within the syndrome measurement circuits from FIG. 3, with probability $p$ [32]. For example, a CNOT is a two gate which has probability $p/15$ a non-trivial two-fold Pauli operator ($I\otimes X,I\otimes Y,\dots,Z\otimes Z$). Often, the noise channel considered for gate errors is the depolarizing channel, but there have also been studies which have considered biased channels so as to reproduce results which may reflect the bias suffered by several experimental qubits.
@@
An interesting approach to the decoding of surface codes under circuit-level noise and several decoding rounds was given in [87], where they consider the errors that may arise from the stabilizing circuit. When a set of errors is indistinguishable (they trigger the same measurement qubits), they are considered equal events under the combination of their probabilities. With these considerations in mind, a Tanner graph can be developed considering all syndrome measurement rounds. A graphic illustration is provided in FIG. 32, in the top figure one can see a planar representation of the stabilizing circuit of a distance-3 surface code. Note that by following the CNOT order given by the violet numbers one can compute all the CNOT gates of the surface code stabilizing circuit with only 4 steps avoiding at all time that one qubit experiences two CNOT gates at the same time. The measurement qubits, which are also defined as detectors in the literature, are denoted with $D_{i}$ labels. After several stabilizing measurements we can describe the overall syndrome through the change in parities of the detectors as shown in FIG. 31, i.e. the syndrome element $D_{i}^{t}$ will be 1 if the detector $D_{i}$ has changed its measurement outcome between the measurement time instances t1t-1 and $t$. Within the top figure, one of the CNOT gates are written in red, in the bottom figure, a Tanner graph relates the possible Pauli errors  [^34] Pauli operator to the left operates on the detector $D_{6}$, while the one on the right operates on the data qubit. which may occur in the noisy CNOT gate with the detector events they would trigger. Consider the case of a $ZX$ error, which is highlighted in orange lines. The $X$ operator on the data qubit which is applied on the circuit time step $3$ will propagate to $D_{3}$ in the following circuit time step following the logic of FIG. 3 and producing a non-trivial $D_{3}^{t}$ due to the parity change in its measurement. In the following round of measurements the $X$ error from the data qubit will propagate to $D_{2}$ changing its parity and producing a non-trivial $D_{2}^{t+1}$. On the other hand, the $Z$ error correspondent to the detector will produce a measurement error, changing its parity and resulting in a non-trivial $D_{6}^{t}$. After another stabilizing round, the parity of the detector changes again since there is no measurement error and so it returns to the previous state, making $D_{6}^{t+1}$ also be non-trivial. FIG. 32 only contains the Tanner graph correspondent to a CNOT error, in order to consider the entire surface code circuit level noise, one should consider all CNOT error Tanner graphs omitting operators which trigger the same set of operators, doing so allows to compute BP, which aids other decoding methods such as MWPM or UF in their task of decoding a syndrome [87]. This method has also been considered for the BPOSD method in [238], where the authors decode a low connection QLDPC code by considering the gate and idling Pauli errors.
@@
Figure 32: In the top, a distance-3 surface code with the CNOT gates from the stabilizing circuit shown. The violet numbers indicate the order at which the CNOT gates are applied within the circuit of every measurement qubit, which themselves are labelled as $D_{i}$. Graphical representation of the circuit-level Tanner graph from a single CNOT, which is denoted in red. The figure is inspired from [87].
@@
Usually, it is considered that a distance-$d$ surface code requires $\mathcal{O}(d)$ syndrome extraction rounds per decoding round [87, 107, 83] in order to preserve the distance property [^35] A measurement error within a check which repeats itself over half the times of syndrome extraction rounds will produce a logical error.. Considering several syndrome extractions further complicates the decoding process, since it significantly increases the number of nodes within decoding graphs and the connectivity of the graph. Specifically, a surface code of arbitrary distance-$d$, which has $d^{2}-1$ checks, results to have $d(d^{2}-1)$ checks. Moreover, even by considering multiple decoding rounds for dealing with faulty measurements, the threshold of the codes decrease dramatically due to such effect caused by circuit-level noise when compared to just data qubit noise [32, 87, 107].
@@
In conclusion, measurement errors pose a significant challenge to the successful implementation of the surface code. The accurate extraction of syndromes is crucial for effective error correction. Faulty-measurements lead to lower code thresholds, implying that the hardware requirements for error correction are more stringent. Therefore, dealing with circuit-level noise and noisy measurements is critical to reach the fault-tolerant quantum computing paradigm.

### v5

@@ 6 Discussion
Following the present discussion, and as it has been seen throughout the text, the decoding stage stands as a pivotal operation within the successful functioning of, not only surface codes, but any correcting code in general. One should be aware of the fact that the named code threshold depends on three aspects: the code family in consideration, the decoding method for the code and the noise model in consideration [68, 69]. In this sense, it is straightforward to see that designing decoding algorithms is crucial in terms of the ability of a family of codes to correct errors. The intrinsic locality and degeneracy of surface codes allows for several decoding processes to present themselves as valid candidates for future real device implementation. Nevertheless, in addition to the code threshold for a decoding method, two other critical aspects stand as limiting factors for the implementation of surface codes decoding in a real-time fashion when implemented on experimental hardware: run time and circuit-level noise.
@@
Decoding runtime refers to the actual time that a decoder requires to ouptut an estimate of the error that has corrupted the logical qubit encoded by a code. Hence, this inherently imposes a delay in the error correcting system that is hazardous for its actual operation. This comes from the fact that, once the syndrome measurement has been done, the data qubits will still continue to suffer from errors as they will not stop to decohere. Thus, if the delay between measuring the checks and applying the recovery operation is too high, then the algorithm will fail, with high probability, the actual error configuration when recovery is executed (See appendix A for a discussion of why real-time decoding is required). This is especially important by taking into account that the decoherence times of state-of-the-art qubits are short and the generation of syndrome data is done a fast rate [^51] Note that this is not true for all qubit technologies such as ion traps, but those present additional problems such as the fact that their operation times, quantum gates, are very slow. [27]. As seen in through this tutorial, it is pretty fair to state that maintaining code threshold while reducing the actual complexity of a decoder is a hard problem. This is a result of the trade-off of being fast against considering more error configurations, which actually relates to code performance. This logic comes from the fact that the traditional comparison of accurate and fast is done in terms of code threshold and worst case complexity. However, it is a recent trend to benchmark runtime by expected complexity, i.e. an average of the number of operation required to decode different error patterns. The rationale behind this is that worst-case events are generally exponentially rare, implying that their impact will not be as high in code performance as other more frequent cases. This effect should be more important whenever hardware noise improves lower to sub-threshold physical error rates and, thus, expected runtime might be a better benchmark for decoding algorithms. In this sense, designing faster decoding algorithms that are able to maintain the capabilities of the code in terms of decoding accuracy is a critical issue for the QEC field. It is important to state that significant advances in this direction are being obtained with proposals such as sparse blossom [44] or Fusion Blossom [96]. Furthermore, recently, sliding-window decoding schemes have been proposed that allow parallel decoding of the syndrome data by means of matching like decoding algorithms (MWPM and UF), thus, dealing with the exponential backlog problem of real-time decoding while maintaining logical fidelity [182, 183].
@@
  • The rectangular rotated code: using a rotated code with rectangular shape makes the minimal weight logical error $Z_{L}$ to require more physical $Z$ operators, increasing the $Z$-distance $d_{Z}$ and, consequently, increasing the number of $Z$-errors it can correct: $(d_{Z}-1)/2$ [106]. On the negative side, this is at the expense of increasing the number of necessary data qubits to $d_{X}d_{Z}$. In FIG. 34, a rectangular code of distances $d_{X}=5,d_{Z}=7$ is shown, notice how the logical operator $Z_{L}$ requires additional Pauli operators than for the $X_{L}$ operator. This code construction may use the reviewed decoders in order to estimate the channel errors.
@@
    Figure 34: Graphical representation of a $d_{X}=5,d_{Z}=7$ rectangular rotated planar code. $Y_{L}=X_{L}Z_{L}$ operator is also represented.
@@
  • Clifford deformation: in Clifford deformed surface codes, the check qubits are conjugated respect to a unitary operator from the Clifford group [185] so that other valid stabilizers are obtained [86]. By means of this stabilizer modifications, the new stabilizer codes can be constructed so that they are more susceptible to $Z$-noise. For example, the XY-code is a surface code which, instead of having $X$ and $Z$-checks, has $Y$ and $X$-checks, both of which are susceptible to $Z$-noise. A graphical depiction of the XY-code is given in the top figure of FIG. 35. As a result of such susceptibility to $Z$-errors, the associated decoder can make use of the correlation between $X$ and $Y$-errors, i.e. $XY=Z$, in order to be more accurate in detecting $Z$-noise [65, 66, 141]. Another popular Clifford deformed surface code is the XZZX surface code [32, 186]. For this code, all check qubits act equally on their surrounding data qubits operating two $X$ and two $Z$ operators as shown in the bottom figure of FIG. 35. One can see that the commutation between the checks is maintained, since adjacent checks either anti-commute twice or directly commute. An interesting characteristic of the XZZX code is that there is only one set of $Z$ physical operators which produce a logical operator. As is shown in FIG. 35, all the products of this logical operator with the stabilizer set produce physical operators which are composed of $X$, $Y$ and $Z$ operators. This feature lowers the probability of a logical phase-flip, $Z_{L}$, error under a highly $Z$-biased noise model. Moreover, for the case in which the $Z$-bias is infinite, e.g. a pure dephasing channel, the XZZX surface code operates as a series of disjoint repetition codes that can be decoded independently. The data qubits of such repetition codes are the top-left to bottom-right diagonals of the overall XZZX code, as can be seen in the bottom of FIG. 35, where the diagonal consisting of two $Z$ physical errors produces syndromes in their top-left and bottom-right sides. Given this condition, the threshold of the XZZX code reaches an outstanding $50\%$ [32, 186], while reducing the amount of required qubits significantly. Therefore, it can be seen that a bias towards $Z$-errors is not only not problematic, but also beneficial if the code and decoder are tailored for such model.
@@
    Figure 35: On the top, a graphical representation of a rotated planar XY code with the operators of the stabilizer generators. We represent $Y$-checks with blue to distinguish them from $Z$-checks (green troughout the text). On the bottom, the same representation but for the rotated planar XZZX code altogether with a $Z_{L}$ operator. We represent all checks with the brown color since all of them are equal now consisting on both $X$ and $Z$ stabilizers.
@@
  • $X/Z$ correlation: another important approach to improve the performance of surface codes is to make decoders aware of the fact that $Y$-errors represent a correlation between the $X$ and $Z$-errors. One downside of the independent $X$ and $Z$ decoding of CSS codes such as CSS surface codes is that it makes the decoding blind of the correlation among those errors, i.e. it is usually considered that $p_{Y}=p_{X}p_{Z}$. However, as seen in Section 4, the Pauli noise models have $p_{X}\approx p_{Y}$ (both in biased and depolarizing) and, thus, by not taking this into account the decoder underestimates the frequency in which $Y$-errors corrupt the data qubits of the code. A method in order to circumvent this problem (without explicitly changing the structure of the code) consists in first decoding for one of the errors and, afterwards, using the partial recovered error in order to decode the other error. E.g., given an error in a CSS surface code, first decode the $X$-error, and then consider that, for recovering for the $Z$-error, every qubit has an updated probability of undergoing a $Z$-error defined by:
@@ Acknowledgements
We want to acknowledge Nicolas Delfosse, Joschka Roffe, Pavel Panteleev, Christopher Chubb, David Tuckett, Michael Newman, Manabu Hagiwara, Iñigo Barasoain and Javier Oliva for fruitful discussions.
@@ Appendix A The backlog problem and why QEC requires real-time decoding
As discussed in the main text, the decoding algorithms for surface codes need to be fast enough to operate in real time. This is necessary because running fault-tolerant algorithms outside the reach of classical simulation requires the use of quantum algorithms that apply $T=\text{diag}(1,e^{i\frac{\pi}{4}})$ gates over the logical space of the code. The use of $T$ gates guarantees that the code supports the universal gate set Clifford+T that can be used to compile any quantum computation, a result known from the Solovay-Kitaev theorem [192]. However, the Eastin-Knill theorem states that no quantum error correction code can implement a universal gate set in a transversal manner  [^52] A transversal logical gate is one that is implemented by performing physical gates over the physical qubits of the code in a manner that avoids the propagation of physical errors throughout the code. For example, the surface codes allows the application of a logical $X$-gate by applying $X$ operators on the physical qubits horizontally from boundary to boundary (Recall FIG. 4).. In this regard, surface codes support transversal implementation of logical Clifford gates but lack a transversal version of the logical $T$-gate. Fortunately, there are ways to circumvent the Eastin-Knill theorem with the purpose of achieving universal fault-tolerant computation. The most popular method is known as magic state injection [182]. A magic state injection protocol can be seen in FIG. 36, where the protocol requires the distillation of a magic state [43] and the injection of such a state to the logical qubit by means of a teleportation circuit.
@@
Coming back to the topic of decoding algorithms, note that if one restricts the operation of the QEC to the set of Clifford gates, then due to the fact that the Clifford group is the normalizer of the Pauli group, it is not necessary to know the logical state for applying the circuit [43]. Since Clifford gates map Paulis into Paulis, the only thing that changes is the interpretation of the final measurements after the computation. Therefore, when running such circuits, one can record the syndrome history during the computation and post-process the errors using the recorded syndrome information. In this sense, one can construct a quantum memory without requiring real-time decoding [43]. However, the scenario drastically changes whenever useful computations are taken into account [^53] By useful we mean quantum computations that cannot be efficiently simulated by classical means and, hence, require $T$-gates.. That is because when injecting the magic state, one must be certain if the $S$-gate after the measurement should be applied or not. Such a decision can only be reliably performed if the syndrome data before the application of said gate is decoded [182]. If the correction of such a Clifford error is ignored, the resulting error might propagate into a much more complicated error that will corrupt the computation [43]. Therefore, in a fault-tolerant quantum computer, the classical processing of the syndrome data, that is, the decoding algorithm; should be fast enough so that syndrome data does not accumulate into an undecodable degree. This accumulation of syndrome data due to the latency of the decoding algorithm is usually referred as the backlog problem [43, 182]. $r_{\mathrm{proc}}$ is defined as the rate at which the decoder processes the syndrome data while $r_{\mathrm{gen}}$ is the rate at which the syndrome data is generated (both quantities in bauds). As discussed by Terhal in [43], if the ratio
$r_{\mathrm{gen}}/r_{\mathrm{proc}}=f>1,$ (45)
then a small initial backlog in the processing of the syndrome data will result in an exponential slow down as a function of the $T$-gate depth of the algorithm. More concretely, Terhal lower bounded the running time of $k$ $T$-depth quantum algorithms as $cf^{k}$ when $f>1$, for some constant $c$ [43, 182]. Thus, the straightforward conclusion is that decoding algorithms must be fast enough to ensure that the injections of the magic states are delayed by only a constant amount of time (instead of an exponentially increasing amount of time).
@@ Appendix B Multiple-round decoding for measurement errors
Throughout this article, surface code decoders have been described as correcting errors arising from data qubit errors, i.e. considering that check measurements are noiseless. However, this simplified perspective is overly optimistic because circuit-level noise arising from gate and SPAM errors make syndrome measurement to be noisy too. As mentioned in the main article section on error models and decoding, circuit level noise will lead to incorrect syndromes. If we attempt to decode an erroneous syndrome, the recovered error will likely result in a logical error. Therefore, a new approach is needed to account this type of errors in the code. The common technique consists in performing multiple syndrome extracting rounds. When a check qubit undergoes a change in its measurement value, it implies a non-trivial syndrome element for such measurement round. Subsequently, decoding algorithms can be employed by connecting check qubits within a particular time instance to themselves in the following time, creating a graph containing space and time information of the errors. Recall that not every decoding algorithm requires this multiple-round measurements as, for example, the cellular automaton decoders [100, 151]. These type of decoders are usually known as single-shot decoders [136]. It is important to say that not every QEC family can admit a single-shot decoding algorithm [136].
Figure 37: Graphical representation of an error in a space-time rotated code. In the figure, each of the depicted rotated planar codes represent one measurement round ($d=3$ in this example). The solid red lines represent the recovered space-time error.
@@
The common approach for considering measurement error noise channels consists in introducing an error probability to the state initialization of the measurement qubit, noisy Haddamard and CNOT gates as described in the error model section and noisy measurement probabilities. Often, this is done by considering all gates within the syndrome measurement circuits from FIG. 3, with probability $p$ [25]. For example, a CNOT is a two-qubit gate which, under probability $p/15$, will produce non-trivial two-fold Pauli operator ($I\otimes X,I\otimes Y,\dots,Z\otimes Z$). Often, the noise channel considered for gate errors is the depolarizing channel, but there have also been studies which have considered biased channels so as to reproduce results which may reflect the bias suffered by several experimental qubits.
@@
An interesting approach to the decoding of surface codes under circuit-level noise and several decoding rounds was given in [65], where they consider the errors that may arise from the stabilizing circuit. When a set of errors is indistinguishable (they trigger the same measurement qubits), they are considered equal events under the combination of their probabilities. With these considerations in mind, a Tanner graph can be developed considering all syndrome measurement rounds. A graphic illustration is provided in FIG. 38, in the top figure one can see a planar representation of the stabilizing circuit of a distance-3 rotated code. Note that by following the CNOT order given by the violet numbers one can compute all the CNOT gates of the rotated code stabilizing circuit with only 4 steps avoiding at all time that one qubit experiences two CNOT gates at the same time. The measurement qubits, which are also defined as detectors in the literature, are denoted with $D_{i}$ labels. After several stabilizing measurements we can describe the overall syndrome through the change in parities of the detectors as shown in FIG. 37, i.e. the syndrome element $D_{i}^{t}$ will be 1 if the detector $D_{i}$ has changed its measurement outcome between the measurement time instances t1t-1 and $t$. Within the top figure, one of the CNOT gates are written in red, in the bottom figure, a Tanner graph relates the possible Pauli errors  [^54] Pauli operator to the left operates on the detector $D_{6}$, while the one on the right operates on the data qubit. which may occur in the noisy CNOT gate with the detector events they would trigger. Consider the case of a $ZX$-error, which is highlighted in orange lines. The $X$ operator on the data qubit which is applied on the circuit time step $3$ will propagate to $D_{3}$ in the following circuit time step following the logic of FIG. 3 and producing a non-trivial $D_{3}^{t}$ due to the parity change in its measurement. In the following round of measurements the $X$-error from the data qubit will propagate to $D_{2}$ changing its parity and producing a non-trivial $D_{2}^{t+1}$. On the other hand, the $Z$-error correspondent to the detector will produce a measurement error, changing its parity and resulting in a non-trivial $D_{6}^{t}$. After another stabilizing round, the parity of the detector changes again since there is no measurement error and so it returns to the previous state, making $D_{6}^{t+1}$ also be non-trivial. FIG. 38 only contains the Tanner graph correspondent to a CNOT error, in order to consider the entire rotated code circuit level noise, one should consider all CNOT error Tanner graphs omitting operators which trigger the same set of operators, doing so allows to compute BP, which aids other decoding methods such as MWPM or UF in their task of decoding a syndrome [65]. This method has also been considered for the BPOSD method in [123], where the authors decode a low connection QLDPC code by considering the gate and idling Pauli errors.
@@
Figure 38: In the top, a distance-3 rotated code with the CNOT gates from the stabilizing circuit shown. The violet numbers indicate the order at which the CNOT gates are applied within the circuit of every measurement qubit, which themselves are labelled as $D_{i}$. Graphical representation of the circuit-level Tanner graph from a single CNOT, which is denoted in red. The figure is inspired from [65].
@@
Usually, it is considered that a distance-$d$ surface code requires $\mathcal{O}(d)$ syndrome extraction rounds per decoding round [61, 65, 87] in order to preserve the distance property [^55] A measurement error within a check which repeats itself over half the times of syndrome extraction rounds will produce a logical error.. Considering several syndrome extractions further complicates the decoding process, since it significantly increases the number of nodes within decoding graphs and the connectivity of the graph. Specifically, a rotated code of arbitrary distance-$d$, which has $d^{2}-1$ checks, results to have $d(d^{2}-1)$ checks. Moreover, even by considering multiple decoding rounds for dealing with faulty measurements, the threshold of the codes decrease dramatically due to such effect caused by circuit-level noise when compared to just data qubit noise [25, 65, 87].
@@
In conclusion, measurement errors pose a significant challenge to the successful implementation of surface codes. The accurate extraction of syndromes is crucial for effective error correction. Faulty-measurements lead to lower code thresholds, implying that the hardware requirements for error correction are more stringent. Therefore, dealing with circuit-level noise and noisy measurements is critical to reach the fault-tolerant quantum computing paradigm.